40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Feb 10 18:00:13 2013, yuta, Summary, LSC, PR2-flipped half-PRC spectra/OLTF POPDCLSCPRM.pngLSCPRCLOLTF.png
    Reply  Mon Feb 11 12:49:54 2013, Jenne, Summary, LSC, Resonant freq change - why? (and passive TT mode freqs) 
Message ID: 8053     Entry time: Sun Feb 10 18:00:13 2013     Reply to this: 8054
Author: yuta 
Type: Summary 
Category: LSC 
Subject: PR2-flipped half-PRC spectra/OLTF 

To compare with future PRMI locking, I measured spectra of POPDC and feedback signal. I also measured openloop transfer function of half-PRC locking.
Beam spot motion was at ~ 2.4 Hz, not 3.3 Hz.

Results:
  Below is uncalibrated spectra of POPDC and LSC feedback signal (C1:LSC-PRM_OUT).
POPDCLSCPRM.png

  Below is openloop transfer function of the half-PRC locking loop. UGF is ~ 120 Hz and phase margin is ~ 45 deg. This agrees with the expected curve.
LSCPRCLOLTF.png

  Data was taken when half-PRC was locked using REFL11_I as error signal and actuating on PRM.


Discussion:

  For comparison, POPDC when PRMI was locked in July 2012: elog #6954 and PRCL openloop transfer function: elog #6950.

  Peak in the spectra of POPDC and feedback signal was at ~ 3.3 Hz in July 2012 PRMI, but it is now at ~ 2.4 Hz in half-PRC. The peak also got broader.
  Is it because of the change in the resonant frequency of the BS-PRM stack? How much the load on BS-PRM changed?
  Or is it because of the change in the resonant frequency of PR2/PR3?

  Phase margin is less now because of gain boost ~ 5 Hz and resonant gain at 24 Hz.
 

ELOG V3.1.3-